log (subscript 4) (2x-15) - log(subscript 4) (x-4)= 1
Answer by ewatrrr(24785)


Hi,


4^1 = (2x-15)/(x-4)
4(x-4) = 2x - 15
4x - 16 = 2x - 15
2x = 1
x = 1/2